OpenStudy (anonymous):

You roll 1 red and 1 white die. What is the probability that the number on the red die is larger than the number on the white die?

OpenStudy (ankit042):

Lets solve this by considering all the cases complete the tabel if white die get number probability values red dice can take 1 1/6 2,3,4,5,6 2 3 4 5 6 complete this table to proceed

OpenStudy (anonymous):



OpenStudy (anonymous):

if the white die rolls 1 then the red die can roll 2,3,4,5,6. probability of getting a number is 1/6

OpenStudy (anonymous):

rolling 1 of white and 2 on red = 1/6 x 1/6

OpenStudy (anonymous):

1/36

OpenStudy (anonymous):

there are 15 cases like these in total therefor 15 x 1/6 x 1/6 = 5/12

OpenStudy (anonymous):

lets count

OpenStudy (anonymous):

i did... using my tree. lol..

OpenStudy (anonymous):

ohhh you divide 3 out of 15/36 to get 5/12 @Mousam

OpenStudy (anonymous):

yup... it is better to simplify fractions to make the final answer look neat...

OpenStudy (anonymous):

argeed :)

OpenStudy (anonymous):

:)

OpenStudy (anonymous):

\((\color{green}1,\color{red}2)\)\((\color{green}1,\color{red}3)\)\((\color{green}1,\color{red}4)\)\((\color{green}1,\color{red}5)\)\((\color{green}1,\color{red}6)\) damn i was having fun coloring....

OpenStudy (anonymous):

haha it looks cool @satellite73

OpenStudy (anonymous):

haha @satellite73

OpenStudy (anonymous):

\((\color{green}2,\color{red}3)\)\((\color{green}2,\color{red}4)\)\((\color{green}2,\color{red}5)\)\((\color{green}2,\color{red}6)\)

OpenStudy (anonymous):

keep going.... you're half way there...

OpenStudy (anonymous):

haha @Mousam

OpenStudy (anonymous):

\[\Huge \begin{array}l\color{red}{\normalsize\text{I}}\color{orange}{\normalsize\text{ }}\color{#9c9a2e}{\normalsize\text{l}}\color{green}{\normalsize\text{o}}\color{blue}{\normalsize\text{v}}\color{purple}{\normalsize\text{e}}\color{purple}{\normalsize\text{ }}\color{red}{\normalsize\text{c}}\color{orange}{\normalsize\text{o}}\color{#9c9a2e}{\normalsize\text{l}}\color{green}{\normalsize\text{o}}\color{blue}{\normalsize\text{r}}\color{purple}{\normalsize\text{}}\end{array}\]